Costs per Equivalent Unit and Production Costs

The following information concerns production in the Forging Department for November. All direct materials are placed into the process at the beginning of production, and conversion costs are incurred evenly throughout the process. The beginning inventory consists of $91,740 of direct materials.

ACCOUNT Work in Process—Forging Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Nov. 1 Bal., 6,600 units, 60% completed 100,056
30 Direct materials, 60,000 units 816,000 916,056
30 Direct labor 56,900 972,956
30 Factory overhead 78,576 ? 1,051,532
30 Goods finished, ? units ?
30 Bal., 5,300 units, 80% completed ?

Cost per equivalent units of $13.60 for Direct Materials and $2.20 for Conversion Costs.

Based on the above data, determine each of the following amounts.

If required, round your interim calculations to two decimal places. Round final answers (a-c) to the nearest dollar.

a. Cost of beginning work in process inventory completed in November.
$

b. Cost of units transferred to the next department during November.
$

c. Cost of ending work in process inventory on November 30.
$

d. Costs per equivalent unit of direct materials and conversion included in the November 1 beginning work in process. If required, round your answers to two decimal places.

Direct materials cost per equivalent unit $
Conversion cost per equivalent unit $

e. The November increase or decrease in costs per equivalent unit for direct materials and conversion from the previous month. If required, round your answers to two decimal places.

Increase or Decrease Amount
Change in direct materials cost per equivalent unit $
Change in conversion cost per equivalent unit $
Answer #1

Unit transferred out = 6600+60000-5300 = 61300

a) Cost of beginning WIP Completed = (6600*40%*2.2)+100056 = 105864

b) Cost of unit completed and transferred out = (15.80*54700)+105864 = 970124

c) Cost of ending WIP = (5300*13.60+4240*2.2)= 81408

d. Costs per equivalent unit of direct materials and conversion included in the November 1 beginning work in process. If required, round your answers to two decimal places.

Direct materials cost per equivalent unit $91740/6600 = 13.90
Conversion cost per equivalent unit $8316/3960 = 2.1

e. The November increase or decrease in costs per equivalent unit for direct materials and conversion from the previous month. If required, round your answers to two decimal places.

Increase or Decrease Amount
Change in direct materials cost per equivalent unit Decrease $0.10
Change in conversion cost per equivalent unit Increase $0.10